Hi guys! I have three questions? That I did & keep getting wrong.
1. Evaluate |-x-2y| for x = -2 and y=3
2. Evaluate the expression (ab)^2 for a = 4 and b=3.
3. Evaluate the formula V=bh/3 for b=9 square inched and h= 32 inches.
____in.^3
Thanks.

Answer by ewatrrr(24785)


Hi
Evaluating by substituting values given:
1. Evaluate |-x-2y| for x = -2 and y=3 |-(-2)- 2*3| = |-4| = 4
2. Evaluate the expression (ab)^2 for a = 4 and b=3. (4*3)^2 = 12^2 = 144
3. Evaluate the formula V=bh/3 for b=9 square inched and h= 32 inches
V = 9in^2*32in/3 = 96in^3